# Arg-LLaDA: Argument Summarization via Large Language Diffusion Models and Sufficiency-Aware Refinement

### v3

@@ Abstract
Argument summarization aims to generate concise, structured representations of complex, multi-perspective debates. While recent work has advanced the identification and clustering of argumentative components, the generation stage remains underexplored. Existing approaches typically rely on single-pass generation, offering limited support for factual correction or structural refinement. To address this gap, we introduce Arg-LLaDA, a novel large language diffusion framework that iteratively improves summaries via sufficiency-guided remasking and regeneration. Our method combines a flexible masking controller with a sufficiency-checking module to identify and revise unsupported, redundant, or incomplete spans—yielding more faithful, concise, and coherent outputs. Empirical results on two benchmark datasets demonstrate that Arg-LLaDA surpasses state-of-the-art baselines in 7 out of 10 automatic evaluation metrics. In addition, human evaluations reveal substantial improvements across core dimensions, coverage, faithfulness, and conciseness,validating the effectiveness of our iterative, sufficiency-aware generation strategy.
@@ 1 Introduction
Argument summarization aiming to distill salient points from multi-perspective debates (known as Key Point Analysis (KPA) (Bar-Haim et al., 2020a, b)) or argumentative texts into concise, faithful, and well-structured summaries (Li et al., 2024a) (Shown in Table 1). This task is particularly important for applications in civic discourse analysis (Vecchi et al., 2021), legal decision support (Habernal et al., 2024), and educational technology (Grljevic et al., 2022), where it is essential to capture the reasoning structures and factual bases behind opposing claims. Current pipeline can be decomposed into two critical stages: (1) identifying and collecting the core argumentative components—namely, claims and their supporting or opposing evidence and (2) generating high-quality summaries based on this structured input van der Meer et al. (2024).
@@
Recent advances in argument summarization have concentrated on what we refer to as Stage 1, which either involves: (i) constructing argumentative components from a fixed, pre-collected corpus; or (ii) retrieving such components dynamically from the open Web. Some studies target the former, automatically identifying argumentative elements—such as claims, premises, and stance—via supervised learning methods (Kapadnis et al., 2021) or prompt-based large language models (LLMs) (Tang et al., 2024a, b; Fukuma et al., 2024). Other work further groups semantically and functionally similar arguments into representative key points (Phan et al., 2021; Reimer et al., 2021), employing techniques such as contrastive learning (Alshomary et al., 2021), graph-based (Li et al., 2024b), and density-based clustering (Li et al., 2023; Khosravani et al., 2024). Meanwhile, another strand of work emphasizes dynamic retrieval of contextual evidence from external sources such as online documents or knowledge bases: models query these sources at inference time to retrieve supporting or elaborating context for argumentative components (Shan and Lu, 2025). This retrieval-augmented generation (RAG) paradigm enables LLMs to ground their outputs in external content, enhancing factuality and currency without retraining the model.
@@
However, Stage 2 remains inadequately investigated. First, most existing approaches rely on end-to-end fine-tuned models (Eden et al., 2023; Bar-Haim et al., 2021) or prompt-based LLMs that generate summaries (Tang et al., 2024b; Altemeyer et al., 2025), providing no inherent mechanism for correction. Once the summary is generated, post-hoc editing becomes cumbersome and unsystematic. This greedy decoding strategy often favors highest-probability tokens at each step, resulting in factual inaccuracies or hallucinations (Andriopoulos and Pouwelse, 2023). Second, unlike text summarization, argument summarization demands conciseness, cover different aspects or facets without redundancy and clear argumentative structure (Li et al., 2025c). Existing models seldom guarantee the preservation of argument roles, such as claim-premise relations, or tightly control thematic redundancy (Li et al., 2024a; Liu et al., 2023; Cheng et al., 2022). Consequently, narrative coherence and structural fidelity are frequently sacrificed, making these methods ill-suited for producing argument summaries that meet the standard.
@@
To address these challenges, we introduce Arg-LLaDA, a novel large language diffusion framework specifically tailored for argument summarization. Rather than generating summaries in a one-shot, greedy fashion, Arg-LLaDA employs an iterative remasking and regeneration strategy designed to incrementally refine outputs through multiple editing cycles. Central to our method is an enhanced span-infilling paradigm (Nie et al., 2025) by with a flexible mask-remask controller, which allows for multiple masking strategies—such as sentence-level masking, thereby enabling precise and context-aware rewriting of deficient summary segments while preserving global coherence. Additionally, we incorporate a sufficiency-checking module, which detects unsupported, redundant, or logically incomplete spans based on the principle of conclusion reconstructability. We validate the effectiveness of our design on the ASE and ArgKP benchmarks. Arg-LLaDA outperforms strong baselines such as Gemini 2.5 and DeepSeek-R1, achieving relative improvements with 7 of 10 metrics best. Human evaluation further confirms these gains, with Arg-LLaDA improving coverage, faithfulness, and conciseness by up to 5.0%. Ablation studies demonstrate that both the iterative refinement and sufficiency diagnosis modules contribute independently to these improvements.
@@
In summary, our key contributions are as follows: (1) Arg-LLaDA, a novel self-refinement model that improves argument summaries via sufficiency-guided remasking and iterative generation. (2) We empirically validate our approach on benchmark datasets, demonstrating significant improvements in content fidelity, conciseness, and argumentative structure over prior baselines.
@@ 2 Related work
Argument Summarization. Traditional approaches often relied on extractive methods, such as clustering-based selection (Misra et al., 2017; Reimers et al., 2019; Ajjour et al., 2019), which group similar arguments without producing fluent summaries. Later work reformulated the task as one of claim generation (Wang and Ling, 2016) or aspect-controlled generation (Schiller et al., 2021), enabling abstractive summaries that better capture the underlying stance. A prominent research line is Key Point Analysis (KPA) (Bar-Haim et al., 2020a, b), which identifies representative key points across a collection of arguments. Follow-up work enhances this pipeline through unsupervised clustering (Alshomary et al., 2021), generation-based key point synthesis (Li et al., 2023), and extractive key point selection with structure-aware scoring (Khosravani et al., 2024). Recent studies explore end-to-end generation models that condition on structured argumentative inputs (Chen et al., 2022) or integrate content selection via reinforcement learning (Roit et al., 2023).
@@
Recent work has increasingly focused on diffusion-based modeling, encompassing both continuous and discrete approaches (Li et al., 2025b). Notably, masked diffusion models, a type of discrete diffusion technique, have emerged as the leading variant in terms of performance. Theoretical studies have provided rigorous foundations for these models and verified their competitiveness with autoregressive architectures at the GPT-2 scale (Shi et al., 2024; Ou et al., 2025; Shao et al., 2024; Li et al., 2025a). Building on this progress, LLaDA (Nie et al., 2025) scaled masked diffusion modeling to 8 billion parameters—becoming the first diffusion-based LLM to rival models like LLaMA 3 (Dubey et al., 2024) in a variety of downstream tasks. Its multimodal variant, LLaDA-V You et al. (2025), outperforms the LLMs baseline across several multimodal benchmarks—indicating potential inherent advantages of the masked-diffusion framework in broader, multimodal contexts.
@@ 3.1 Problem Definition
3.2 Diffusion Generation Framework
Inspired by recent advances in masked diffusion language models (Nie et al., 2025), we adopt a non-autoregressive, iterative generation paradigm for argument summarization. Unlike autoregressive decoders, which generate outputs token-by-token in a fixed order and struggle with revising earlier errors, our diffusion-based approach refines summaries through repeated masked denoising steps. This iterative design allows for flexible editing, span-level abstraction, and improved controllability—characteristics well-aligned with the requirements of factual and concise argument summarization.
Training Objective.
Following Nie et al. (2025), the model is trained via a denoising score matching objective. Specifically, given a reference summary $\hat{S}$, we randomly corrupt a subset of tokens $\mathcal{M}\subset\hat{S}$ and optimize the model to reconstruct the original tokens conditioned on the unmasked context and input arguments:
3.3 Inference-Time Sufficiency Guided Masking
Token-level low-confidence remasking. At each inference step, the model produces probability distributions over candidate tokens. We compute token confidence as the maximum probability assigned to any candidate token. Tokens with confidence below a threshold $\tau$ are deemed unreliable, and the lowest-$r$ proportion of such tokens are selected for resampling. This mechanism ensures that local corrections are directed toward genuinely uncertain regions while avoiding unnecessary revisions of fluent and stable tokens. Following prior work on diffusion-based generation (Nie et al., 2025), we apply confidence-based remasking only at inference, not training. This separation ensures that the diffusion backbone learns a broad, task-agnostic denoising capability, while inference-time confidence signals specialize revision to context-specific uncertainties.
@@ 3.5 Training and Implementation Details
We initialize the diffusion model with the LLaDA checkpoint (Nie et al., 2025), pretrained on large-scale web corpora with masked diffusion objectives. Its architecture and noise scheduling are retained. For fine-tuning, we supervise the model on domain-specific argument summarization data. Given input tuples of claims and evidence with reference summaries, we apply stochastic masking (30% ratio). The model is then optimized to denoise the masked summary using the diffusion objective (Section 3.2).
@@
We train a RoBERTa-base encoder (Liu et al., 2019) to predict span-level sufficiency labels. Each training instance consists of a sentence-level span $s$, its associated claim $c_{i}$, and supporting evidence $E_{i}$. Labels are automatically generated through controlled perturbations, including hallucinations, contradictions, and omissions. The classifier is trained with binary cross-entropy loss and applied only during inference to guide selective span masking.
@@ 4 Experimental Setup
Broadly speaking, we aim to investigate the effectiveness of our proposed Arg-LLaDA framework in addressing the limitations of current argument summarization methods. Specifically, we pose the following research questions: (i) Does Arg-LLaDA improve the overall quality of argument summaries compared to standard one-pass or prompt-based methods? (ii) To what extent does the sufficiency-based span diagnosis contribute to reducing redundancy and enhancing faithfulness? (iii) What is the impact of different sufficiency checking modules?
@@
We compare Arg-LLaDA against ten state-of-the-art baselines, categorized into three groups: (i) Zero-shot proprietary LLMs, including GPT-4 OpenAI (2023), GPT-o3  [^1] https://openai.com/index/introducing-o3-and-o4-mini/, Gemini 2.5 (Anil et al., 2023), Claude 3.7  [^2] https://www.anthropic.com/news/claude-3-7-sonnet, and DeepSeek-R1 (DeepSeek-AI et al., 2025), which are closed-source commercial models accessed via API without any task-specific adaptation; (ii) Zero-shot open-source LLMs, such as LLaMA 3 (Dubey et al., 2024), LLaDA (Nie et al., 2025), Mistral (Jiang et al., 2023), Gemma (Rivière et al., 2024) and FlanT5 (Chung et al., 2024), which are publicly released and evaluated in their instruction-tuned form without additional training; (iii) Full-shot open-source LLMs, which are further fine-tuned on benchmark datasets to adapt them explicitly to the argument summarization task.
@@
The ArgKP dataset consists of approximately 27k labeled argument–key point pairs covering 31 diverse debate topics, with each pair annotated for stance polarity (Bar-Haim et al., 2020a, b). Following prior work, we partition the dataset into training and test sets in a 28:3 ratio while preserving topic distribution. In addition, we employ and extand the ASE dataset, which provides a comprehensive end-to-end benchmark for argument (Li et al., 2024a) summarization. It spans the full pipeline—from claim and evidence detection to the generation and evaluation of debate-ready summaries—enabling holistic assessment of argument generation systems. For detail about dataset process see Appendix A.
@@
Table 2: Comparison of Arg-LLaDA against zero-shot and fine-tuned LLMs on two argument summarization datasets. We report ROUGE (R-1, R-2, R-L), BLEURT, and BERTScore. The best and second-best results are respectively bolded and underlined.
@@
We report ROUGE (Lin, 2004), which common used in text summarization. While ROUGE captures surface-level similarity, it may overlook semantic equivalence in cases of lexical variation. To do so, we consider BERTScore (Zhang et al., 2020) and BLEURT (Sellam et al., 2020) evaluates summary quality based on contextualized token embeddings obtained from a pre-trained BERT model. They better account for semantic similarity than ROUGE and are less sensitive to exact word matching.
@@
To address these shortcomings, we complement our automatic evaluation with a human assessment focused on three essential dimensions: Coverage, Faithfulness, and Conciseness (Altemeyer et al., 2025). These criteria reflect core desiderata of argumentative summaries—namely, the ability to accurately capture all salient perspectives, avoid unsupported or fabricated content, and maintain clarity without redundancy.
@@
All outputs are anonymized and randomly ordered to mitigate evaluator bias. We report the average scores across all annotators, and assess inter-annotator agreement via Krippendorff’s alpha. This evaluation offers a more comprehensive view of summary quality, complementing automatic scores with expert human judgment.

### v4

@@ Abstract
Argument summarization aims to generate concise, structured representations of complex, multi-perspective debates. While recent work has advanced the identification and clustering of argumentative components, the generation stage remains underexplored. Existing approaches typically rely on single-pass generation, offering limited support for factual correction or structural refinement. To address this gap, we introduce Arg‑LLaDA, a novel large language diffusion framework that iteratively improves summaries via sufficiency-guided remasking and regeneration. Our method combines a flexible masking controller with a sufficiency-checking module to identify and revise unsupported, redundant, or incomplete spans—yielding more faithful, concise, and coherent outputs. Empirical results on two benchmark datasets demonstrate that Arg‑LLaDA surpasses state-of-the-art baselines in 7 out of 10 automatic evaluation metrics. In addition, human evaluations reveal substantial improvements across core dimensions, coverage, faithfulness, and conciseness,validating the effectiveness of our iterative, sufficiency-aware generation strategy.
@@ 1 Introduction
Argument summarization aiming to distill salient points from multi-perspective debates (known as Key Point Analysis (KPA) (DBLP:conf/acl/Bar-HaimEFKLS20; DBLP:conf/emnlp/Bar-HaimKEFLS20)) or argumentative texts into concise, faithful, and well-structured summaries (DBLP:conf/acl/0074WSBMZZWHLN24) (Shown in Table 1). This task is particularly important for applications in civic discourse analysis (DBLP:conf/acl/VecchiFJL20), legal decision support (DBLP:journals/ail/HabernalFRBGDB24), and educational technology (DBLP:journals/eis/GrljevicBK22), where it is essential to capture the reasoning structures and factual bases behind opposing claims. Current pipeline can be decomposed into two critical stages: (1) identifying and collecting the core argumentative components—namely, claims and their supporting or opposing evidence and (2) generating high-quality summaries based on this structured input DBLP:conf/eacl/MeerVJM24.
@@
Recent advances in argument summarization have concentrated on what we refer to as Stage 1, which either involves: (i) constructing argumentative components from a fixed, pre‑collected corpus; or (ii) retrieving such components dynamically from the open Web. Some studies target the former, automatically identifying argumentative elements—such as claims, premises, and stance—via supervised learning methods (DBLP:conf/argmining/KapadnisPPMN21) or prompt‑based large language models (LLMs) (DBLP:conf/eacl/TangZD24; DBLP:conf/acl/Tang0DC24; DBLP:journals/corr/abs-2404-18371). Other work further groups semantically and functionally similar arguments into representative key points (DBLP:conf/argmining/PhanNND21; DBLP:conf/argmining/ReimerLHA21), employing techniques such as contrastive learning (DBLP:conf/argmining/AlshomaryGSHSCP21), graph‑based (DBLP:conf/naacl/LiJHXCH24), and density‑based clustering (DBLP:conf/acl/LiSBN23; DBLP:conf/naacl/Khosravani0T24). Meanwhile, another strand of work emphasizes dynamic retrieval of contextual evidence from external sources such as online documents or knowledge bases: models query these sources at inference time to retrieve supporting or elaborating context for argumentative components (DBLP:journals/ijon/ShanL25). This retrieval‑augmented generation (RAG) paradigm enables LLMs to ground their outputs in external content, enhancing factuality and currency without continually training the model.
@@
However, Stage 2 remains inadequately investigated. Firstly, most existing approaches rely on end-to-end fine-tuned models (eden2023welcome; DBLP:conf/acl/Bar-HaimEKFS20) or prompt-based LLMs that generate summaries (DBLP:conf/acl/Tang0DC24; DBLP:journals/corr/abs-2503-00847) in an auto-regressive manner, providing no inherent mechanism for correction. Once the summary is generated, post-hoc editing becomes cumbersome and unsystematic. This greedy decoding strategy often favors highest-probability tokens at each step, resulting in factual inaccuracies or hallucinations (DBLP:journals/corr/abs-2309-16459). Secondly, unlike text summarization, argument summarization demands conciseness, covers different aspects or facets without redundancy, with a clear argumentative structure (li2025large). Existing models rarely guarantee the preservation of argument roles, such as claim-premise relations, or tightly control thematic redundancy (DBLP:conf/acl/0074WSBMZZWHLN24; DBLP:journals/corr/abs-2304-07666; DBLP:conf/acl/ChengBHYZS22). As a result, these methods often compromise narrative coherence and structural fidelity, rendering them ill-suited for producing argument summaries that meet the required standards.
@@
To address these challenges, we introduce Arg‑LLaDA, a novel large language diffusion framework specifically tailored for argument summarization. Rather than generating summaries in a one-shot, greedy fashion, Arg‑LLaDA employs an iterative remasking and regeneration strategy designed to incrementally refine outputs through multiple editing cycles. Central to our method is a novel enhanced span-infilling paradigm (DBLP:journals/corr/abs-2502-09992) with a flexible mask-remask controller, which allows for multiple masking strategies—such as sentence-level masking, thereby enabling precise and context-aware rewriting of deficient summary segments while preserving global coherence. Additionally, we incorporate a sufficiency-checking module, which detects unsupported, redundant, or logically incomplete spans based on the principle of conclusion reconstructability. We validate the effectiveness of our design on the ASE and ArgKP benchmarks. Arg‑LLaDA outperforms state-of-the-art LLM baselines such as Gemini 2.5 and DeepSeek-R1, achieving relative improvements with 7 of 10 metrics best. Human evaluation further validates these gains, with Arg‑LLaDA improving coverage, faithfulness, and conciseness by up to 5.0% compared to baselines. Ablation studies demonstrate that the iterative refinement and sufficiency diagnosis modules contribute orthogonally to model improvements, showing their independent effectiveness.
@@
In summary, our key contributions are as follows: (1) Arg‑LLaDA, a novel self-refinement model that improves argument summaries via sufficiency-guided remasking and iterative generation. (2) We empirically validate our approach on benchmark datasets, demonstrating significant improvements in content fidelity, conciseness, and argumentative structure over prior baselines.
@@ 2 Related work
Argument Summarization. Traditional approaches often relied on extractive methods, such as clustering-based selection (DBLP:journals/corr/abs-1711-00092; DBLP:conf/acl/ReimersSBDSG19; DBLP:conf/emnlp/AjjourAWS19), which group similar arguments without producing fluent summaries. Later work reformulated the task as one of claim generation (DBLP:conf/naacl/WangL16) or aspect-controlled generation (DBLP:conf/naacl/SchillerDG21), enabling abstractive summaries that better capture the underlying stance. A prominent research line is Key Point Analysis (KPA) (DBLP:conf/acl/Bar-HaimEFKLS20; DBLP:conf/emnlp/Bar-HaimKEFLS20), which identifies representative key points across a collection of arguments. Follow-up work enhances this pipeline through unsupervised clustering (DBLP:conf/argmining/AlshomaryGSHSCP21), generation-based key point synthesis (DBLP:conf/acl/LiSBN23), and extractive key point selection with structure-aware scoring (DBLP:conf/naacl/Khosravani0T24). Recent studies explore end-to-end generation models that condition on structured argumentative inputs (DBLP:conf/nips/ChenL0022) or integrate content selection via reinforcement learning (DBLP:conf/acl/RoitFSACDGGHKMG23).
@@
Recent work has increasingly focused on diffusion-based modeling, encompassing both continuous and discrete approaches (DBLP:journals/corr/abs-2503-02445). Notably, masked diffusion models, a type of discrete diffusion technique, have emerged as the leading variant in terms of performance. Theoretical studies have provided rigorous foundations for these models and verified their competitiveness with autoregressive architectures at the GPT-2 scale (DBLP:conf/nips/ShiHWDT24; DBLP:conf/iclr/OuNXZSLL25; DBLP:journals/corr/abs-2402-03300; DBLP:journals/corr/abs-2506-07584). Building on this progress, LLaDA (DBLP:journals/corr/abs-2502-09992) scaled masked diffusion modeling to 8 billion parameters—becoming the first diffusion-based LLM to rival models like LLaMA 3 (DBLP:journals/corr/abs-2407-21783) in a variety of downstream tasks. Its multimodal variant, LLaDA‑V you2025llada, outperforms the LLMs baseline across several multimodal benchmarks—indicating potential inherent advantages of the masked-diffusion framework in broader, multimodal contexts.
@@ 3.2 Sufficiency Diagnosis Module
3.3 Diffusion Generation Framework
Inspired by recent advances in masked diffusion language models (DBLP:journals/corr/abs-2502-09992), we adopt a non-autoregressive, iterative generation paradigm for argument summarization. Unlike autoregressive decoders, which generate outputs token-by-token in a fixed order and struggle with revising earlier errors, our diffusion-based approach refines summaries through repeated masked denoising steps. This iterative design allows for flexible editing, span-level abstraction, and improved controllability—characteristics well-aligned with the requirements of factual and concise argument summarization.
Training Objective.
Following DBLP:journals/corr/abs-2502-09992, the model is trained via a denoising score matching objective. Specifically, given a reference summary $\hat{S}$, we randomly corrupt a subset of tokens $\mathcal{M}\subset\hat{S}$ and optimize the model to reconstruct the original tokens conditioned on the unmasked context and input arguments:
3.4 Inference-Time Sufficiency Guided Masking
Token-level low-confidence remasking. At each inference step, the model produces probability distributions over candidate tokens. We compute token confidence as the maximum probability assigned to any candidate token. Tokens with confidence below a threshold $\tau$ are deemed unreliable, and the lowest-$r$ proportion of such tokens are selected for resampling. This mechanism ensures that local corrections are directed toward genuinely uncertain regions while avoiding unnecessary revisions of fluent and stable tokens. Following prior work on diffusion-based generation (DBLP:journals/corr/abs-2502-09992), we apply confidence-based remasking only at inference, not training. This separation ensures that the diffusion backbone learns a broad, task-agnostic denoising capability, while inference-time confidence signals specialize revision to context-specific uncertainties.
@@ 3.5 Training and Implementation Details
We initialize the diffusion model with the LLaDA checkpoint (DBLP:journals/corr/abs-2502-09992), pretrained on large-scale web corpora with masked diffusion objectives. Its architecture and noise scheduling are retained. For fine-tuning, we supervise the model on domain-specific argument summarization data. Given input tuples of claims and evidence with reference summaries, we apply stochastic masking (30% ratio). The model is then optimized to denoise the masked summary using the diffusion objective (Section 3.3).
@@
We train a RoBERTa-base encoder (DBLP:journals/corr/abs-1907-11692) to predict span-level sufficiency labels. Each training instance consists of a sentence-level span $s$, its associated claim $c_{i}$, and supporting evidence $E_{i}$. Labels are automatically generated through controlled perturbations, including hallucinations, contradictions, and omissions. The classifier is trained with binary cross-entropy loss and applied only during inference to guide selective span masking.
@@ 4 Experimental Setup
Broadly speaking, we aim to investigate the effectiveness of our proposed Arg‑LLaDA framework in addressing the limitations of current argument summarization methods. Specifically, we pose the following research questions: (i) Does Arg‑LLaDA improve the overall quality of argument summaries compared to standard one-pass or prompt-based methods? (ii) To what extent does the sufficiency-based span diagnosis contribute to reducing redundancy and enhancing faithfulness? (iii) What is the impact of different sufficiency checking modules?
@@
We compare Arg‑LLaDA against ten state-of-the-art baselines, categorized into three groups: (i) Zero-shot proprietary LLMs, including GPT-4 DBLP:journals/corr/abs-2303-08774, GPT-o3  [^1] https://openai.com/index/introducing-o3-and-o4-mini/, Gemini 2.5 (DBLP:journals/corr/abs-2312-11805), Claude 3.7  [^2] https://www.anthropic.com/news/claude-3-7-sonnet, and DeepSeek-R1 (DBLP:journals/corr/abs-2501-12948), which are closed-source commercial models accessed via API without any task-specific adaptation; (ii) Zero-shot open-source LLMs, such as LLaMA 3 (DBLP:journals/corr/abs-2407-21783), LLaDA (DBLP:journals/corr/abs-2502-09992), Mistral (DBLP:journals/corr/abs-2310-06825), Gemma (DBLP:journals/corr/abs-2408-00118) and FlanT5 (DBLP:journals/jmlr/ChungHLZTFL00BW24), which are publicly released and evaluated in their instruction-tuned form without additional training; (iii) Full-shot open-source LLMs, which are further fine-tuned on benchmark datasets to adapt them explicitly to the argument summarization task.
@@
The ArgKP dataset consists of approximately 27k labeled argument–key point pairs covering 31 diverse debate topics, with each pair annotated for stance polarity (DBLP:conf/acl/Bar-HaimEFKLS20; DBLP:conf/emnlp/Bar-HaimKEFLS20). Following prior work, we partition the dataset into training and test sets in a 28:3 ratio while preserving topic distribution. In addition, we employ and extand the ASE dataset, which provides a comprehensive end-to-end benchmark for argument (DBLP:conf/acl/0074WSBMZZWHLN24) summarization. It spans the full pipeline—from claim and evidence detection to the generation and evaluation of debate-ready summaries—enabling holistic assessment of argument generation systems.
@@
Table 2: Comparison of Arg‑LLaDA against zero-shot and fine-tuned LLMs on two argument summarization datasets. We report ROUGE (R-1, R-2, R-L), BLEURT, and BERTScore. The best and second-best results are respectively bolded and underlined.
@@
We report ROUGE (lin2004rouge), which common used in text summarization. While ROUGE captures surface-level similarity, it may overlook semantic equivalence in cases of lexical variation. To do so, we consider BERTScore (DBLP:conf/iclr/ZhangKWWA20) and BLEURT (DBLP:conf/acl/SellamDP20) evaluates summary quality based on contextualized token embeddings obtained from a pre-trained BERT model. They better account for semantic similarity than ROUGE and are less sensitive to exact word matching.
@@
To address these shortcomings, we complement our automatic evaluation with a human assessment focused on three essential dimensions: Coverage, Faithfulness, and Conciseness (DBLP:journals/corr/abs-2503-00847). These criteria reflect core desiderata of argumentative summaries—namely, the ability to accurately capture all salient perspectives, avoid unsupported or fabricated content, and maintain clarity without redundancy.
@@
All outputs are anonymized and randomly ordered to mitigate evaluator bias. We report the average scores across all annotators, and assess inter-annotator agreement via Krippendorff’s alpha. This evaluation offers a more comprehensive view of summary quality, complementing automatic scores with expert human judgment. Detail could be find at Appendix B.